OpenStudy (anonymous):

G(x) = 3x2 - 2x - 1 G(6) =

OpenStudy (aakashsudhakar):

G(6) = 3(6)^2 - 2(6) - 1 G(6) = 3(36) - 12 - 1 G(6) = 108 - 12 - 1 G(6) = 95

OpenStudy (anonymous):

G(6) is the value of function G when x = 6 so, G96) = 3(6^2) -2(6) -1 = 3(36) - 12 -1 = 108 -12 -1 = 95